OpenStudy (anonymous):

I've asked this question on brainly.com and yahoo answers but haven't got an answer yet. I'll give a medal for help! Isaac's family drove 100 miles to Yakima on Friday and then they drove back home on Sunday. On the way to Yakima their average speed was 10 mph faster than on their return trip. a) If we represent their rate on their way to Yakima by x, write an expression to show the time spent driving to Yakima. b) If we represent their rate on their way to Yakima by x, write an expression to show the time spent driving home from Yakima. c) Write a simplified expression to represent how much longer it took them to drive home.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

having a little bit of trouble with this one... time = distance / rate (or speed) distance (d) = 100 rate = x.... so expression for time going by the formula would be t = 100/x now that driving TO Yakira driving home ... d = 100 rate = well...since going up there is 10 mph faster....then I think you could just subtract 10 d = 100/(x-10)...but this is where I kinda get messed up on.....lets get somebody here to help

OpenStudy (johnweldon1993):

That is what I was thinking as well :)

OpenStudy (texaschic101):

or would the first one...going to...be 100/(x + 10) and coming back be 100/x

OpenStudy (johnweldon1993):

Nope I believe you have it right above...since they specify that "x" is the rate going there...so that would need to be the first part

OpenStudy (texaschic101):

oh...ok...that makes sense

OpenStudy (anonymous):

Thx so much @texaschic101 but how would I have to figure out c? Would the simplified expression be like a combination of both expressions in A-B?

OpenStudy (texaschic101):

I am sorry but I have to go now...lost track of time..sorry

OpenStudy (johnweldon1993):

The only thing is you have the equations as \[\large D = \frac{100}{x}\] and \[\large D = \frac{100}{(x - 10)}\] Those D's should be T's \[\large T_{there} = \frac{100}{x}\] \[\large T_{home} = \frac{100}{(x - 10)}\]

OpenStudy (anonymous):

oh okay thank you though @texaschic101

OpenStudy (johnweldon1993):

And then yes I would assume you just subtract the two

OpenStudy (anonymous):

I would subtract 2 from both expressions?

OpenStudy (johnweldon1993):

No, you would subtract the two equations... Like Time it took to go there - Time it took to come home

OpenStudy (anonymous):

Oh okay thank you @johnweldon1993

OpenStudy (johnweldon1993):

Of course :)